Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

What is the probability that the sum will equal 9??

Expert replies
by gmatpup » Mon Nov 21, 2011 4:24 am
A = {2, 3, 4, 5}

B= {4, 5, 6, 7, 8}

Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. WHat is the probability that the sum of the two integers will equal 9?

A. 0.15
B. 0.20
C. 0.25
D. 0.30
E. 0.33

Answer is B


Can someone please tell me how to approach questions such as this? Thanks so much!! :)
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Mon Nov 21, 2011 4:43 am
gmatpup wrote:A = {2, 3, 4, 5}

B= {4, 5, 6, 7, 8}

Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. WHat is the probability that the sum of the two integers will equal 9?

A. 0.15
B. 0.20
C. 0.25
D. 0.30
E. 0.33

Answer is B


Can someone please tell me how to approach questions such as this? Thanks so much!! :)
There are total 20 ways to sum.
4 of the sums equals 9.
so, the required probability = 4/20 = 1/5 = .20

The correct answer is B.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by chaitanyabhansali » Tue Nov 22, 2011 11:41 am
Anurag@Gurome wrote:
gmatpup wrote:A = {2, 3, 4, 5}

B= {4, 5, 6, 7, 8}

Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. WHat is the probability that the sum of the two integers will equal 9?

A. 0.15
B. 0.20
C. 0.25
D. 0.30
E. 0.33

Answer is B


Can someone please tell me how to approach questions such as this? Thanks so much!! :)
No. of ways of selecting one number from set A = 4C1 = 4
No. of ways of selecting one number from set B = 5C1 = 5

Total Ways = 4 * 5 = 20

There are 4 ways in which the sum is 9
1) 2(from A) + 7(from B) = 9
2) 3(from A) + 6(from B) = 9
3) 4(from A) + 5(from B) = 9
4) 5(from A) + 4(from B) = 9

Therefore, required probabiity = 4/20 = 0.2
IMO: B
Join the discussion

by tuanquang269 » Tue Nov 22, 2011 8:29 pm
4 x 5 = 20 ways to choose couple of number each from set A and B

There are 4 couple have sum of 9 include: (2,7) (3,6) (4,5) (5,4)

Probability = 4/20 = 20%
Join the discussion

by MBA.Aspirant » Wed Nov 23, 2011 1:15 am
gmatpup wrote:A = {2, 3, 4, 5}

B= {4, 5, 6, 7, 8}


Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. WHat is the probability that the sum of the two integers will equal 9?

A. 0.15
B. 0.20
C. 0.25
D. 0.30
E. 0.33

Answer is B


Can someone please tell me how to approach questions such as this? Thanks so much!! :)
Total events = 4 * 5= 20

favorable events = 4 (2+7), (5+4), (4+5), (3+6)

probability = 4/20 = 0.2 (B)
Join the discussion

by sk8legend408 » Wed Nov 23, 2011 5:28 am
To make it even more simpler:

Each number in set A can match exactly one number in set B to equal 9.

So if you choose any number from set A, the probability that you choose the right number from set B is 1/5, because only one number out of the 5 will be the right match to equal 9. Therefore the answer is .20.
Join the discussion

hi

by Jeff@TargetTestPrep » Mon Apr 09, 2018 3:37 pm
gmatpup wrote:A = {2, 3, 4, 5}

B= {4, 5, 6, 7, 8}

Two integers will be randomly selected from the sets above, one integer from set A and one integer from set B. WHat is the probability that the sum of the two integers will equal 9?

A. 0.15
B. 0.20
C. 0.25
D. 0.30
E. 0.33
Let’s first determine the total number of outcomes. We have 4 numbers in set A, and 5 in set B, and since we are selecting 1 number from each set, the total number of outcomes is 4 x 5 = 20.

For our favorable outcomes, we need to determine the number of ways we can get a number from set A and a number from set B to sum to 9. We are selecting from the following two sets:

A = {2, 3, 4, 5}

B = {4, 5, 6, 7, 8}

We will denote the first number as from set A and the second from set B. Here are the pairings that yield a sum of 9:

2,7
3,6
4,5
5,4

We see that there are 4 favorable outcomes. Thus, the probability is 4/20 = 1/5 = 0.2.

Answer: B

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion